The demand for tickets to an Ethiopian camparada film is given by D(p)=20,000-10,000p, where p is the price of tickets. If the price of tickets is 12 birr. Calculate price elasticity of demand for tickets and draw the demand curve?

1
Expert's answer
2021-11-09T13:25:23-0500

The demand for tickets to an Ethiopian camparada film is given by D(p)=200,000-10,000p, where p is the price of tickets. If the price of tickets is 12 birr. Calculate price elasticity of demand for tickets and draw the demand curve?


Given :

D(p)=200,000-10,000 p

Change in demand/change in price = -10,000

At price 12 Birr change demand units will be 

D(p)= 200,000 - 10,000*12

= 80,000 units

Elasticity in demand= Change in demand/ Change in price unit

= -10,000

Given :

D(p)=200,000 -10,000 p

Change in demand/change in price = -10,000

At price 12 Birr change demand units will be 

D(p)= 200,000- 10,000*12

= 80,000 units

Elasticity in demand= Change in demand/ Change in price unit

"= -10000 \times \frac{12}{80000} \\

= -1.5(approximately)"

Therefore price elasticity of demand is -1.5

Price elasticity is a measure that refers to change in the % of quantity demanded due to % change in price.

As the price elasticity is negative it implies that as the price of the ticket increases demand will keep on falling.